Solve the following linear systems by substitution
m=8j
-m+2=-7j

Answer by jim_thompson5910(35256)

-m+2=-7j

-8j+2=-7j

2=-7j+8j

2=j

j = 2

m = 8j

m = 8(2)

m = 16